# Low-Light Video Enhancement via Spatial-Temporal Consistent Illumination and Reflection Decomposition

### v1

@@ 3.2 Spatial-Temporal Consistent Decomposition
Motivation. Video can be conceptualized as static/dynamic 3D multi-view data Wang et al. (2023b, a). Prior research has demonstrated that the multi-view data itself can be harnessed to decompose the view-dependent and view-independent elements, which respectively encapsulate the intrinsic texture and view-altering illuminations Wang et al. (2023a). Thus, once we establish correspondence relationships among $I_{d,t}$ for all $t$, we can apply the view-independent constraint to derive the reflectance component, denoted as $\hat{R}_{n,t}$, as it represents the intrinsic properties of the target scene. Subsequently, obtaining the illumination component, denoted as $\hat{L}_{n,t}$, becomes achievable through the utilization of the inherent image reconstruction loss, in conjunction with adherence to the continuity assumption.
Implementation of correspondences. Obtaining correspondences for $I_{d,t}$, $t\in[1,T]$, can be a challenging task due to the presence of various degradations, including visibility issues and noise. Fortunately, we have access to corresponding normal-light data, which significantly aids in establishing these correspondences, as illustrated in Fig. 2, where $I_{d,t}$ and $I_{n,t}$ are pixel-wise aligned.
Given two frames $I_{n,t_{1}}$ and $I_{n,t_{2}}$, correspondences can be determined using the prediction network $F$ Edstedt et al. (2023). We assume the detection of $M$ correspondences, denoted as ${c_{m}=(x_{1,m},y_{1,m},x_{2,m},y_{2,m}),m\in[1,M]}$. Each correspondence, represented by $c_{m}$, consists of four coordinates: $(x_{1,m},y_{1,m})$ represents the pixel coordinate in $I_{n,t_{1}}$, while $(x_{2,m},y_{2,m})$ signifies the coordinate in $I_{n,t_{2}}$. Furthermore, each correspondence is associated with an uncertainty value $u_{m}$, determined through a data-driven approach.
Constraints formulation. To satisfy the constraint that the corresponding pixels in the reflectance part are matched, we can formulate $\mathcal{L}(R_{t})$ as
$\small\mathcal{L}(R_{t_{1},t_{2}})=\mathbb{E}_{m\in[1,M]}u_{m}\|\hat{R}_{n,t_{1}}[x_{1,m},y_{1,m}]-\hat{R}_{n,t_{2}}[x_{2,m},y_{2,m}]\|.$ (5)
The reconstruction item in Eq. 3 can be denoted as
$\small\mathcal{L}_{R,t}=\|I_{n,t}-\hat{L}_{n,t}\otimes\hat{R}_{n,t}\|.$ (6)
Although the illumination part can be obtained via Eqs. 5 and 6, we further incorporate a continuity constraint to align with real-world properties. For a pixel in the video frame, denoted as $p$, we introduce a spatial continuity loss on the predicted $\hat{L}_{n,t}$ as
$\small\mathcal{L}(L_{t})=\mathbb{E}_{t}[v_{t}^{p}\times[\partial_{x}\hat{L}_{n,t}(p)]^{2}+u_{t}^{p}\times[\partial_{y}\hat{L}_{n,t}(p)]^{2}],$ (7)
@@
$\small v_{t}^{p}=(\|\partial_{x}\mathcal{T}_{t}(p)\|^{1.2}+\Delta)^{-1},u_{t}^{p}=(\|\partial_{y}\mathcal{T}_{t}(p)\|^{1.2}+\Delta)^{-1},$ (8)
where $\mathcal{T}_{t}$ represents the logarithmic transformation of $I_{d,t}$, and $\Delta$ is a small constant (set to 0.0001) used to avoid division by zero. The diagram for the decomposition implementation is shown in Fig. 2.
@@ 3.3 Dual Network for Consistent Decomposition
Overview and motivation. To enhance the consistency of decomposition, it’s crucial for various frames within a video to mutually share features. By incorporating shared features and simultaneous supervisions, the features used for synthesizing Retinex parameters across different frames can exhibit greater consistency. When the video is enhanced frame by frame without feature propagation, the decomposition results tend to be suboptimal due to the inherently challenging nature of maintaining consistency. This observation is empirically validated in our ablation study. Current LLVE methods primarily leverage multiple-frame inputs to facilitate propagation Wang et al. (2021a); Fu et al. (2023a). However, these strategies come with a notable cost in terms of aligning features, potentially affecting inference efficiency.
In this paper, we introduce an efficient strategy for propagation using a dual approach. This approach entails loading two frames, specifically target frame $I_{d,t_{1}}$ and reference frame $I_{d,t_{2}}$ (where $I_{d,t_{2}}$ can be randomly selected from the temporal neighbors of $I_{d,t_{1}}$ during training, and is set as the closest frame of $I_{d,t_{1}}$ during inference), propagating features at the deepest layer of the network through our cross-frame interaction (as depicted in Fig. 3), and concurrently supervising these dual outputs to guarantee consistency.
Figure 4: Visual samples: derive the Retinex decomposition results of enhanced normal-light outputs from the input low-light video data on different datasets.
Feature propagation via long-range attention and short-range fusion in CFIM. Suppose the feature of $I_{d,t_{1}}$ and $I_{d,t_{2}}$ are denoted as $f_{d,t_{1}}$ and $f_{d,t_{2}}$ which is extracted from the same encoder $E$ in the network $f$. To complete the propagation, we initially employ a long-range cross-frame attention operation in the feature space. A traditional attention operation Vaswani et al. (2017) typically consists of the query vector $q$, the key vector $k$, and the value vector $v$. The attention relationship is established using $A(q,k,v)=softmax(q\times k^{T})\times v$, where $\times$ represents matrix multiplication. To enable cross-frame attention, we process the feature via dual paths, as
$\hat{f}_{d,t_{1}}=A(f_{d,t_{1}},f_{d,t_{1}},f_{d,t_{1}})+A(f_{d,t_{1}},f_{d,t_{2}},f_{d,t_{2}}),$ (9)
$\hat{f}_{d,t_{2}}=A(f_{d,t_{2}},f_{d,t_{2}},f_{d,t_{2}})+A(f_{d,t_{2}},f_{d,t_{1}},f_{d,t_{1}}).$
@@
$\small\bar{f}_{d,t_{1}},\bar{f}_{d,t_{2}}=S(C(\hat{f}_{d,t_{1}}\oplus\hat{f}_{d,t_{2}})),$ (10)
where $\oplus$ represents channel concatenation, $C$ refers to the convolution network, and $S$ signifies channel dimension splitting. The decomposition results for $I_{d,t_{1}}$ and $I_{d,t_{2}}$ are obtained by processing $\bar{f}_{d,t_{1}}$ and $\bar{f}_{d,t_{2}}$ through the decoder. We have confirmed that our dual network with synchronous supervision for the outputs of $I_{d,t_{1}}$ and $I_{d,t_{2}}$ can yield superior decomposition and enhancement results compared to individual enhancement strategies.
3.4 Overall Loss Function
During the training, we find it is better to adopt the dual training strategy, i.e., for each $I_{d,t_{1}}$, we sample its neighboring reference $I_{d,t_{2}}$ and constrain their outputs simultaneously. Thus, the loss function can be written as
$\small\mathcal{L}_{v}=\mathbb{E}_{t_{1},t_{2}}[\mathcal{L}_{R,t_{1}}+\mathcal{L}_{R,t_{2}}+\lambda_{1}(\mathcal{L}(L_{t_{1}})+\mathcal{L}(L_{t_{2}}))+\lambda_{2}\mathcal{L}(R_{t_{1},t_{2}})],$ (11)
where $\lambda_{1}$ and $\lambda_{2}$ are the loss weights, and each loss term is defined in Eqs. 5, 6 and 7. Moreover, we observe that the performance of our framework is robust to the choice of the reference frame.
@@ 4 Experiments
4.1 Datasets
Our evaluation is conducted on four publicly available datasets, which encompass a wide range of real-world videos with diverse motion patterns and degradations. (1) SMID Chen et al. (2019): This dataset consists of static videos with ground truths obtained through long exposure. We followed the procedure provided by SMID, converting the videos from the RAW domain to the sRGB domain using rawpy’s default ISP. (2) SDSD Wang et al. (2021a): SDSD is a dynamic video dataset collected through electromechanical equipment, including both indoor and outdoor subsets. (3) DID Fu et al. (2023a): DID is another dynamic video dataset characterized by pronounced camera motion, strict spatial alignment, and diverse scene content. In comparison to SDSD, DID offers more data and is a multi-illuminance, multi-camera LLVE dataset. We follow DID’s division into a training set, a test set, and a validation set, maintaining a ratio of 3:1:1. (4) DAVIS Pont-Tuset et al. (2017): DAVIS is a dataset that includes both scene dynamics and camera motions. To simulate low-light degradation in DAVIS, we followed the methodology outlined in Zhang et al. (2021), further incorporating noise degradation.
4.2 Implementation Details
We conducted experiments on all datasets using the same network structure. All modules were trained end-to-end, with the learning rate initialized at $4e^{-4}$ for all layers, adapted by a cosine learning scheduler. The batch size used was 4. Correspondences were computed using the SOTA method DKM Edstedt et al. (2023). Additionally, no data augmentation was applied during training. All experiments were executed on a single NVIDIA A100 GPU.
@@ 4.2 Implementation Details
SNR Xu et al. (2022) 27.30 0.84 23.23 0.82 28.49 0.81 24.85 0.90
SMG Xu et al. (2023b) 27.82 0.86 25.17 0.81 28.03 0.78 25.14 0.85
PairLLE Fu et al. (2023b) 23.48 0.71 20.04 0.65 22.70 0.63 22.56 0.82
RetinexFormer Cai et al. (2023) 26.56 0.79 22.80 0.77 29.15 0.82 25.40 0.89
MBLLEN Lv et al. (2018) 22.17 0.66 21.41 0.63 22.67 0.68 24.22 0.86
SMID Chen et al. (2019) 24.84 0.72 23.30 0.67 24.78 0.72 22.28 0.84
SMOID Jiang and Zheng (2019) 24.63 0.70 22.25 0.68 23.64 0.71 22.13 0.85
SDSDNet Wang et al. (2021a) 26.81 0.75 23.08 0.71 26.03 0.75 22.52 0.81
DP3DF Xu et al. (2023a) 27.63 0.74 23.85 0.73 27.19 0.76 22.39 0.88
StableLLVE Zhang et al. (2021) 25.32 0.70 22.47 0.65 24.37 0.69 21.64 0.80
LLVE-SEG Liu et al. (2023) 26.19 0.77 24.09 0.72 25.31 0.74 23.85 0.82
LLVE-CFA Chhirolya et al. (2022) 24.28 0.81 22.64 0.76 24.72 0.70 22.53 0.87
@@
Table 2: Quantitative comparison on the DAVIS dataset.
Methods SNR SMG PairLLE RetinexFormer MBLLEN SMID
PSNR 20.69 20.18 19.57 21.79 18.63 20.51
SSIM 0.710 0.672 0.667 0.723 0.619 0.686
Methods SMOID SDSDNet DP3DF StableLLVE LLVE-SEG Ours
PSNR 20.66 21.22 22.04 21.48 21.45 23.38
SSIM 0.697 0.716 0.749 0.732 0.715 0.782
@@ 4.3 Comparison and Baselines
We conducted a comprehensive comparison with SOTA LLVE methods, including MBLLEN Lv et al. (2018), SMID Chen et al. (2019), SMOID Jiang and Zheng (2019), SDSDNet Wang et al. (2021a), DP3DF Xu et al. (2023a), StableLLVE Zhang et al. (2021), LLVE-SEG Liu et al. (2023), and LLVE-CFA Chhirolya et al. (2022) We also compared our method with SOTA LLIE methods for individual frames, including SNR Xu et al. (2022), SMG Xu et al. (2023b), PairLLE Fu et al. (2023b), RetinexFormer Cai et al. (2023). All these methods were trained on each dataset using their respective released code and hyper-parameters (e.g., the training epoch that is set to guarantee the convergence). Note that all these baselines are trained on our unified data split for a fair comparison. The splits slightly differ from those in baselines’ original papers, so the scores of the baselines may vary from the original papers.
Furthermore, Table 2 provides a summary of the comparative results on DAVIS. In comparison to the degradation synthesis strategy as presented in Zhang et al. (2021), we have extended our approach to include the degradation of a low-light noise term. As a result, our synthesized data encompasses dynamic scenes with pronounced invisibility and perturbations, posing a considerable challenge. As illustrated in Table 2, our method consistently yields the highest PSNR and SSIM values, reaffirming the effectiveness of our approach.
Qualitative result. Besides the quantitative comparison, we present visual comparisons with the selected baselines. Fig. 5 showcases the visual comparisons on SDSD and SMID, while Fig. 6 displays visual cases from DID and DAVIS. In general, the results enhanced by our approach exhibit a more natural appearance, including accurate color, well-balanced brightness, enhanced contrast, and precise details without noise. Furthermore, our framework’s results show fewer artifacts in regions with complex textures, and they appear cleaner and sharper compared to the results produced by other methods. This distinction is particularly notable, as most baselines perform well in simpler areas.
@@
Input
RetinexFormer
SDSDNet
DP3DF
Ours
GT
Figure 5: Visual comparisons on SDSD-indoor (top), SDSD-outdoor (middle), and SMID (bottom). The results of our proposed framework, i.e., “Ours”, demonstrate better visual perception with less noise, clearer visibility, and more enhanced details.
Input
RetinexFormer
SDSDNet
DP3DF
Ours
GT
Input
RetinexFormer
SDSDNet
DP3DF
Ours
GT
Figure 6: Comparisons on DID (top) and DAVIS (bottom). “Ours” has less noise, better visibility and details.
4.4 Ablation Study
We assess the critical components of our framework through five ablation cases. (1) “w/o LR” that removes the constraint on learning the reflectance part. (2) “w/o LL” where the constraint on learning the illumination part is removed. (3) “w/o C.F.”: we eliminate the cross-frame attention and fusion in the network, resulting in each frame being enhanced individually. (4) “with M.I.”: This refers to using multiple neighboring frames as input and employing the temporal alignment of deformable convolution. (5) “w/o Dual”: the dual learning strategy is removed, meaning that the loss is applied to only one frame in each iteration.
The results are summarized in Table 3. By comparing “w/o LR” with the full setting (“Full”), we can clearly demonstrate the effectiveness of our proposed constraint for the reflectance map. Similarly, the significant advantage of “Full” over “w/o LL” highlights the significance of the proposed illumination continuity constraint. Furthermore, we can validate the importance of mutually propagating features among different frames to achieve spatial-temporal consistent decomposition by comparing “w/o C.F.” with “Full”. Our proposed propagation strategy via the dual network structure proves to be more effective than the common method of using multi-frame inputs for propagation in videos, as seen in the comparison between “w/o M.I.” and “Full”. The dual learning setting, where we simultaneously supervise the two outputs of the dual network, also proves to be valuable, as evidenced by the comparison between “w/o Dual” and “Full”.
Table 3: Ablation study on SDSD, SMID, and DID.

### v2

@@ 3.2 Spatial-Temporal Consistent Decomposition
Motivation. Video can be conceptualized as static/dynamic 3D multi-view data (Wang et al., 2023b, a). Prior research has demonstrated that the multi-view data itself can be harnessed to decompose the view-dependent and view-independent elements, which respectively encapsulate the intrinsic texture and view-altering illuminations (Wang et al., 2023a). Thus, once we establish correspondence relationships among $\boldsymbol{I}_{d,t}$ for all $t$, we can apply the view-independent constraint to derive $\hat{\boldsymbol{R}}_{n,t}$, as it represents the intrinsic properties of the target scene. Subsequently, obtaining $\hat{\boldsymbol{L}}_{n,t}$ becomes achievable through the utilization of the inherent image reconstruction loss, in conjunction with adherence to the continuity assumption.
Implementation of Correspondences. Obtaining correspondences for $\boldsymbol{I}_{d,t}$, $t\in[1,T]$, can be a challenging task due to the presence of various degradations, including visibility issues and noise. Fortunately, we have access to corresponding normal-light data, which significantly aids in establishing these correspondences, as illustrated in Fig. 2, where $\boldsymbol{I}_{d,t}$ and $\boldsymbol{I}_{n,t}$ are pixel-wise aligned.
Given two frames $\boldsymbol{I}_{n,t_{1}}$ and $\boldsymbol{I}_{n,t_{2}}$, correspondences can be determined using the prediction network $F$ (Edstedt et al., 2023). We assume the detection of $M$ correspondences, denoted as ${c_{m}=(x_{1,m},y_{1,m},x_{2,m},y_{2,m}),m\in[1,M]}$. Each correspondence, represented by $c_{m}$, consists of four coordinates: $(x_{1,m},y_{1,m})$ represents the pixel coordinate in $\boldsymbol{I}_{n,t_{1}}$, while $(x_{2,m},y_{2,m})$ signifies the coordinate in $\boldsymbol{I}_{n,t_{2}}$. Furthermore, each correspondence is associated with an uncertainty value $u_{m}$, determined through a data-driven approach.
Figure 3: The lightweight cross-frame interaction mechanism (CFIM) propagates different frames’ features, along with cross-frame attention and spatial-channel fusion. Cross-frame interaction can be employed in the deep feature space of arbitrary single-image encoder-decoder frameworks, and we choose U-Net here.
Constraints Formulation. To satisfy the constraint that the corresponding pixels in $\boldsymbol{R}_{t}$ are matched, we have $\mathcal{O}(\boldsymbol{R}_{t})$ as
$\small\mathcal{O}(\boldsymbol{R}_{t_{1},t_{2}})=\mathop{\mathbb{E}}\limits_{m\in[1,M]}u_{m}\|\hat{\boldsymbol{R}}_{n,t_{1}}[x_{1,m},y_{1,m}]-\hat{\boldsymbol{R}}_{n,t_{2}}[x_{2,m},y_{2,m}]\|.$ (5)
The reconstruction item in Equation 3 can be denoted as
$\mathcal{O}_{rec,t}=\|\boldsymbol{I}_{n,t}-\hat{\boldsymbol{L}}_{n,t}\otimes\hat{\boldsymbol{R}}_{n,t}\|.$ (6)
Although the view-dependent part can be obtained via the reconstruction constraint, We further incorporate a continuity constraint to align with real-world properties. For a pixel in the video frame, denoted as $p$, we introduce a spatial continuity objective on the predicted $\hat{\boldsymbol{L}}_{n,t}$ as the following loss function
$\mathcal{O}(\boldsymbol{L}_{t})=\mathbb{E}_{t}[v_{t}^{p}\times[\partial_{x}\hat{\boldsymbol{L}}_{n,t}(p)]^{2}+u_{t}^{p}\times[\partial_{y}\hat{\boldsymbol{L}}_{n,t}(p)]^{2}],$ (7)
@@
$\small v_{t}^{p}=(\|\partial_{x}\boldsymbol{U}_{t}(p)\|+\Delta)^{-1},u_{t}^{p}=(\|\partial_{y}\boldsymbol{U}_{t}(p)\|+\Delta)^{-1},$ (8)
where $\boldsymbol{U}_{t}$ represents the logarithmic transformation of $\boldsymbol{I}_{d,t}$, and $\Delta$ is a small constant (set to 0.0001) used to avoid division by zero. The diagram is shown in Fig. 2.
@@ 3.3 Dual Network for Consistent Decomposition
Overview and Motivation. To enhance the consistency of decomposition, it’s crucial for various frames within a video to mutually share features. By incorporating shared features and simultaneous supervision, the features used for synthesizing decomposition parameters across different frames can exhibit greater consistency. When the video is enhanced frame by frame without feature propagation, the decomposition results tend to be suboptimal due to the inherently challenging nature of maintaining consistency. This observation is empirically validated in our ablation study. Current LLVE methods primarily leverage multiple-frame inputs to facilitate propagation (Wang et al., 2021; Fu et al., 2023a). However, these strategies come with a notable cost in terms of aligning features, potentially affecting inference efficiency.
In this paper, we introduce an efficient strategy for propagation using a dual approach. This approach entails loading two frames, specifically target frame $\boldsymbol{I}_{d,t_{1}}$ and reference frame $\boldsymbol{I}_{d,t_{2}}$ (where $\boldsymbol{I}_{d,t_{2}}$ can be randomly selected from the temporal neighbors of $\boldsymbol{I}_{d,t_{1}}$ during training, and is set as the closest frame of $\boldsymbol{I}_{d,t_{1}}$ during inference), propagating features at the deepest layer of the network through our cross-frame interaction (as depicted in Fig. 3), and concurrently supervising these dual outputs to guarantee consistency.
Feature Propagation via Spatial-varying Fusion. Suppose the feature of $\boldsymbol{I}_{d,t_{1}}$ and $\boldsymbol{I}_{d,t_{2}}$ are denoted as $\boldsymbol{f}_{d,t_{1}}$ and $\boldsymbol{f}_{d,t_{2}}$ which is extracted from the same encoder $M_{E}$ in the network $f$. To complete the propagation, we initially employ a long-range cross-frame attention operation in the feature space. A traditional attention operation (Vaswani et al., 2017) typically consists of the query vector $\boldsymbol{Q}$, the key vector $\boldsymbol{K}$, and the value vector $\boldsymbol{V}$. The attention relationship is established using $A(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})=\text{softmax}(\boldsymbol{Q}\times\boldsymbol{K}^{\top})\times\boldsymbol{V}$, where $\times$ represents matrix multiplication. To enable cross-frame attention, we process the feature via dual paths, as
$\boldsymbol{f}^{\prime}_{d,t_{1}}=A(\boldsymbol{f}_{d,t_{1}},\boldsymbol{f}_{d,t_{1}},\boldsymbol{f}_{d,t_{1}}),\boldsymbol{f}^{\prime\prime}_{d,t_{1}}=A(\boldsymbol{f}_{d,t_{1}},\boldsymbol{f}_{d,t_{2}},\boldsymbol{f}_{d,t_{2}}),$ (9)
$\boldsymbol{f}^{\prime}_{d,t_{2}}=A(\boldsymbol{f}_{d,t_{2}},\boldsymbol{f}_{d,t_{2}},\boldsymbol{f}_{d,t_{2}}),\boldsymbol{f}^{\prime\prime}_{d,t_{2}}=A(\boldsymbol{f}_{d,t_{2}},\boldsymbol{f}_{d,t_{1}},\boldsymbol{f}_{d,t_{1}}).$
@@
$(\displaystyle\hat{\boldsymbol{f}}_{d,t_{1}},\hat{\boldsymbol{f}}_{d,t_{2}}=S(C($$(\boldsymbol{f}_{d,t_{1}}+\boldsymbol{f}^{\prime}_{d,t_{1}}+\boldsymbol{f}^{\prime\prime}_{d,t_{1}})\oplus$ (10)
$(\boldsymbol{f}_{d,t_{2}}+\boldsymbol{f}^{\prime}_{d,t_{2}}+\boldsymbol{f}^{\prime\prime}_{d,t_{2}}))),$
where $\oplus$ represents channel concatenation, $C$ refers to the convolution network, and $S$ signifies channel dimension splitting. The decomposition results for $\boldsymbol{I}_{d,t_{1}}$ and $\boldsymbol{I}_{d,t_{2}}$ are obtained by processing $\hat{\boldsymbol{f}}_{d,t_{1}}$ and $\hat{\boldsymbol{f}}_{d,t_{2}}$ through the decoder $M_{D}$. We have confirmed that our dual network with synchronous supervision for the outputs of $\boldsymbol{I}_{d,t_{1}}$ and $\boldsymbol{I}_{d,t_{2}}$ can yield superior decomposition and enhancement results compared to individual enhancement strategies.
The Role of the Cross-attention in CFIM. The cross-attention mechanism facilitates information propagation across different frames, aligning with our training procedure that requires spatial consistency for $\boldsymbol{R}$. Moreover, during training, CFIM brings varying filtered features from randomly selected reference images ($\boldsymbol{I}_{d,t_{2}}$) into the backbone (i.e., the part without feature propagation), where the input is the image at the current time step $\boldsymbol{I}_{d,t_{1}}$. This setup enables CFIM to guide the backbone in learning how to adaptively leverage diverse knowledge from reference images, resulting in robust and effective feature spaces within the backbone, without the need for expensive temporal alignment. In other words, the backbone performs strongly without a reference image during inference (though, of course, having a suitable reference image to provide additional useful information generally improves performance), as confirmed by the ablation study. Additionally, CFIM can leverage useful cross-frame information when it complements the features in the backbone, further enhancing the decomposition process.
3.4 Overall Objective to Compute Loss Function
During the training, we find it is better to adopt the dual training strategy, i.e., for each $\boldsymbol{I}_{d,t_{1}}$, we sample its neighboring reference $\boldsymbol{I}_{d,t_{2}}$ and constrain their outputs simultaneously. Thus, the loss function can be written as
$\mathcal{O}_{v}=\mathbb{E}_{t_{1},t_{2}}[$ $\mathcal{O}_{rec,t_{1}}+\mathcal{O}_{rec,t_{2}}+$ (11)
$\lambda_{1}(\mathcal{O}(\boldsymbol{L}_{t_{1}})+\mathcal{O}(\boldsymbol{L}_{t_{2}}))+\lambda_{2}\mathcal{O}(\boldsymbol{R}_{t_{1},t_{2}})],$
where $\lambda_{1}$ and $\lambda_{2}$ are the loss weights, and each loss term is defined in Equation 5, Equation 6 and Equation 7.
@@ 4 Experiments
SNR 27.30 0.84 23.23 0.82 28.49 0.81 24.85 0.90
SMG 27.82 0.86 25.17 0.81 28.03 0.78 25.14 0.85
PairLLE 23.48 0.71 20.04 0.65 22.70 0.63 22.56 0.82
RetinexFormer 26.56 0.79 22.80 0.77 29.15 0.82 25.40 0.89
MBLLEN 22.17 0.66 21.41 0.63 22.67 0.68 24.22 0.86
SMID 24.84 0.72 23.30 0.67 24.78 0.72 22.28 0.84
SMOID 24.63 0.70 22.25 0.68 23.64 0.71 22.13 0.85
SDSDNet 26.81 0.75 23.08 0.71 26.03 0.75 22.52 0.81
DP3DF 27.63 0.74 23.85 0.73 27.19 0.76 22.39 0.88
StableLLVE 25.32 0.70 22.47 0.65 24.37 0.69 21.64 0.80
LLVE-SEG 26.19 0.77 24.09 0.72 25.31 0.74 23.85 0.82
LLVE-CFA 24.28 0.81 22.64 0.76 24.72 0.70 22.53 0.87
BLLRVE 26.02 0.74 23.50 0.71 26.15 0.73 23.25 0.81
VRGSS 27.81 0.76 24.28 0.75 27.84 0.78 24.51 0.87
ZRLLE-PQP 25.53 0.71 22.42 0.73 26.36 0.77 22.84 0.80
JDAE 26.19 0.74 23.37 0.76 26.45 0.76 23.53 0.82
@@
4.1 Datasets
Our evaluation is conducted on four publicly available datasets, which encompass a wide range of real-world videos with diverse motion patterns and degradations, including SMID (Chen et al., 2019), SDSD (Wang et al., 2021), DID (Fu et al., 2023a), and DAVIS (Pont-Tuset et al., 2017).
4.2 Implementation Details
Experimental Details. We conducted experiments on all datasets using the same network structure. $T$ is set as 5 in the experiment. All modules were trained end-to-end, with the learning rate initialized at $4e^{-4}$ for all layers, adapted by a cosine learning scheduler. The batch size used was 4. Correspondences were computed using the SOTA method DKM (Edstedt et al., 2023). We used the pre-trained weights of DKM during training because it is designed for general indoor and outdoor scenes. Its generalization ability is supported by its extensive training data and state-of-the-art training strategy, as demonstrated in the original DKM paper and subsequent studies (Zhu and Liu, 2023; Edstedt et al., 2024) (through evaluations on unseen real-world indoor and outdoor scenes).
@@ 4.3 Comparison and Baselines
We conducted a comprehensive comparison with SOTA LLVE methods, including MBLLEN (Lv et al., 2018), SMID (Chen et al., 2019), SMOID (Jiang and Zheng, 2019), SDSDNet (Wang et al., 2021), DP3DF (Xu et al., 2023a), StableLLVE (Zhang et al., 2021), LLVE-SEG (Liu et al., 2023), LLVE-CFA (Chhirolya et al., 2022), BLLRVE (Zhang et al., 2024), VRGSS (Li et al., 2023), ZRLLE-PQP (Wang et al., 2024), and JDAE (Shi et al., 2024). We also compared our method with SOTA LLIE methods for individual frames, including SNR (Xu et al., 2022), SMG (Xu et al., 2023b), PairLLE (Fu et al., 2023b), RetinexFormer (Cai et al., 2023). All these methods were trained on each dataset using their respective released code and hyper-parameters (e.g., the training epoch that is set to guarantee convergence). Note that all baselines are trained on our unified data split for a fair comparison. The splits slightly differ from those in baselines’ original papers, so the scores of baselines may vary from the original papers.
@@
Figure 4: Visual comparisons on SDSD-indoor (top two rows), SDSD-outdoor (middle two rows), and SMID (bottom two rows). The results of our proposed framework, i.e., “Ours”, demonstrate better visual perception with clearer visibility and more enhanced details.
@@
Figure 5: Comparisons on DID (top two rows) and DAVIS (bottom two rows). “Ours” has better visibility and details.
SNR SMG PairLLE RetinexFormer MBLLEN SMID
PSNR 20.69 20.18 19.57 21.79 18.63 20.51
SSIM 0.710 0.672 0.667 0.723 0.619 0.686
SMOID SDSDNet DP3DF StableLLVE LLVE-SEG Ours
PSNR 20.66 21.22 22.04 21.48 21.45 23.38
SSIM 0.697 0.716 0.749 0.732 0.715 0.782
Table 2: Quantitative comparison on the DAVIS dataset.
Furthermore, Table 2 provides a summary of the comparative results on DAVIS. In comparison to the degradation synthesis strategy as presented in (Zhang et al., 2021), we have extended our approach to include the degradation of a low-light noise term. As a result, our synthesized data encompasses dynamic scenes with pronounced invisibility and perturbations, posing a considerable challenge. As illustrated in Table 2, our method consistently yields the highest PSNR and SSIM values, reaffirming the effectiveness of our approach.
DAVIS SDSD-Indoor SDSD-Outdoor DID
Methods Short Long Short Long Short Long Short Long
SNR 0.027 0.070 0.017 0.060 0.022 0.046 0.025 0.068
RetinexFormer 0.029 0.072 0.018 0.061 0.024 0.043 0.024 0.065
SDSDNet 0.024 0.068 0.012 0.053 0.011 0.038 0.017 0.059
DP3DF 0.021 0.065 0.011 0.050 0.013 0.036 0.019 0.062
StableLLVE 0.023 0.067 0.016 0.055 0.019 0.042 0.023 0.066
LLVE-SEG 0.025 0.071 0.014 0.057 0.016 0.040 0.021 0.064
Ours-R 0.017 0.058 0.010 0.048 0.007 0.030 0.011 0.049
Ours 0.020 0.063 0.012 0.051 0.010 0.034 0.015 0.054
Table 3: The quantitative comparison in terms of short-term (“Short”) and long-term (“Long”) temporal loss. “Ours-R” means the view-independent term produced by ours.
Evaluation for Temporal Consistency. For video processing, the performance of temporal consistency and stability should be evaluated. Thus, we employ the short-term and long-term temporal loss proposed in (Lai et al., 2018) for such temporal evaluation on different datasets. The wrapping operations among frames are computed on the frames with the normal light. Moreover, the long-term loss is computed every 10 frames. The results are shown in Table 3 (we normalize the frame values into [0, 1]). Obviously, our results have lower temporal loss than the baselines, e.g., higher temporal consistency and stability. Specifically, we assess the temporal consistency of the view-independent term. The results presented in Table 3 further validate the constancy.
Qualitative Result. Besides the quantitative comparison, we present visual comparisons with the selected baselines. Fig. 4 showcases the visual comparisons of SDSD and SMID, while Fig. 5 displays visual cases from DID and DAVIS. In general, the results enhanced by our approach exhibit a more natural appearance, including accurate color, well-balanced brightness, enhanced contrast, and precise details. Our results show fewer artifacts in regions with complex textures, and they appear cleaner and sharper compared to results produced by other methods. This distinction is particularly notable, as most baselines perform well in simpler areas.
@@
Table 4: Ablation study on SDSD, SMID, and DID.
4.4 Ablation Study
We assess the critical components of our framework through five ablation cases. (1) “w/o LR” that removes the constraint on learning the view-independent part. (2) “w/o LL” where the constraint on learning the view-dependent part is removed. (3) “w/o C.F.”: we eliminate the cross-frame attention and fusion in the network, resulting in each frame being enhanced individually. (4) “with M.I.”: This refers to using multiple neighboring frames as input and employing the temporal alignment of deformable convolution. (5) “w/o Dual”: the dual learning strategy is removed, meaning that the loss is applied to only one frame in each iteration.
The results are summarized in Table 4. By comparing “w/o LR” with the full setting (“Full”), we can clearly demonstrate the effectiveness of our proposed constraint for the view-independent part. Similarly, the significant advantage of “Full” over “w/o LL” highlights the significance of the proposed continuity constraint. Furthermore, we can validate the importance of mutually propagating features among different frames to achieve spatial-temporal consistent decomposition by comparing “w/o C.F.” with “Full”. Our proposed propagation strategy via the dual network structure proves to be more effective than the common method of using multi-frame inputs for propagation in videos, as seen in the comparison between “w/o M.I.” and “Full”. The dual learning setting, where we simultaneously supervise the two outputs of the dual network, also proves to be valuable, as evidenced by the comparison between “w/o Dual” and “Full”.